OpenStudy (anonymous):

Given the function f(x) = −2x + 8, find x if f(x) = 14

OpenStudy (anonymous):

14 = -2x +8 2x = -6 x = -3

OpenStudy (anonymous):

petite is wrong

OpenStudy (anonymous):

the function f(x) = -2(14)+8 = -20, so f(x) = -20 . (I did this last chapter in precalculus)

OpenStudy (anonymous):

did you read the question?

OpenStudy (anonymous):

Just replace the X with 14 in the equation and replace f(x) so it's y instead.

jimthompson5910 (jim_thompson5910):

no petite is the only one who is correct actually

OpenStudy (anonymous):

it says f(x) is 14, find x ,not x is is 14

OpenStudy (anonymous):

OH

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

my bad, ignore me, i misread the question. im always trying to help but im always wrong, i might as well give up now :)

OpenStudy (anonymous):

\[f(14)=-2(14)+8\] \[f(14)=-28+8\] \[f(14)=-20\]

jimthompson5910 (jim_thompson5910):

you have it backwards overachiever

OpenStudy (amistre64):

f(x) = 14 ; not f(14) ....

jimthompson5910 (jim_thompson5910):

just follow petite's work and answer and you'll be fine op

OpenStudy (anonymous):

oh sorry i read it wrong too

OpenStudy (amistre64):

f(x) = -2x + 8 | 14 = -2x + 8 ; is the correct way to solve it :)

OpenStudy (amistre64):

fenwick, being wrong is fine :) as long as we learn from our mistakes

jimthompson5910 (jim_thompson5910):

sometimes that's the best way to learn